# Personalized Enhanced Federated Multi-View Clustering via Heat-Kernel Tensor Decomposition

### v2

@@ 6.5.1 Global Objective with Personalization
$J_{\text{global}}=\sum_{\ell=1}^{M}\left[\omega_{\ell}J_{FedHK}^{(\ell)}(U_{[\ell]},A_{[\ell]},V_{[\ell]})+\gamma\mathcal{R}_{person}(A_{[\ell]},A_{\text{global}})+\eta\mathcal{R}_{view}(V_{[\ell]},V_{\text{global}}\right]$ (86)
where:
  • $\omega_{\ell}$ represents client-specific weights based on data quality and quantity
  • $\mathcal{R}_{person}(A_{[\ell]},A_{\text{global}}=\sum_{h,k}\|A_{[\ell]}^{h}-A_{\text{global}}^{h}\|_{F}^{2}$ ensures coordination between local and global centers
  • $\mathcal{R}_{view}(V_{[\ell]},V_{\text{global}})=\|V_{[\ell]}-V_{\text{global}}\|_{2}^{2}$ maintains view weight consistency
  • $\gamma,\eta$ control personalization vs. coordination trade-offs
@@ 6.5.3 Personalized Update Mechanism
$A_{\ell}^{h,(t+1)}$ $=\lambda_{\ell}A_{\ell}^{h,(t)}+(1-\lambda_{\ell})A_{\text{global}}^{h,(t+1)}$ (93)
$V_{\ell}^{(t+1)}$ $=\rho_{\ell}V_{\ell}^{(t)}+(1-\rho_{\ell})V_{\text{global}}^{(t+1)}$ (94)
@@
$\lambda_{\ell}^{(t+1)}=\lambda_{\ell}^{(t)}-\beta\nabla_{\lambda_{\ell}}J_{[\ell]}(\lambda_{\ell}A_{\ell}^{(t)}+(1-\lambda_{\ell})A_{\text{global}}^{(t)})$ (95)
6.6 Tensorized Federated Aggregation with Personalization
6.6.1 Tensorized Global Objective with Personalization
$J_{\text{tensor(global)}}=\sum_{\ell=1}^{M}\left[\omega_{\ell}J_{[\ell]}^{tensor}(\mathcal{U}_{[\ell]},\mathcal{G}_{[\ell]},\mathbf{P}_{[\ell]},\mathbf{Q}_{[\ell]},\mathbf{R}_{[\ell]},\mathbf{V}_{[\ell]})+\gamma\mathcal{R}_{tensor}(\mathcal{G}_{[\ell]},\mathbf{P}_{[\ell]},\mathbf{Q}_{[\ell]},\mathbf{R}_{[\ell]})+\eta\mathcal{R}_{view}(V_{[\ell]},V_{\text{global}})\right]$ (96)
where the tensorized regularization terms are:
$\mathcal{R}_{tensor}$ $=\|\mathcal{G}_{[\ell]}-\mathcal{G}_{\text{global}}\|_{F}^{2}+\|\mathbf{P}_{[\ell]}-\mathbf{P}_{\text{global}}\|_{F}^{2}+\|\mathbf{Q}_{[\ell]}-\mathbf{Q}_{\text{global}}\|_{F}^{2}+\|\mathbf{R}_{[\ell]}-\mathbf{R}_{\text{global}}\|_{F}^{2}$ (97)
$\mathcal{R}_{view}(V_{[\ell]},V_{\text{global}})$ $=\|V_{[\ell]}-V_{\text{global}}\|_{2}^{2}$ (98)
@@ 6.6.3 Tensorized Personalized Update Mechanism
Algorithm 4 provides a comprehensive framework for implementing personalized FedHK-PARAFAC2. In order to achieve a personalized FedHK-Tucker, two changes must be made. Initially, adjustments are made to the tensor factor implication stage and its localization tensorized HKC. Secondly, modifications are made to the tensor aggregation stage through the implementation of Tucker decomposition.
Algorithm 4 Personalized FedHK-PARAFAC2
Server Initialization:
 Initialize global tensor factors $\mathcal{G}_{\text{global}}^{(0)},\mathbf{P}_{\text{global}}^{(0)},\mathbf{Q}_{\text{global}}^{(0)},\mathbf{R}_{\text{global}}^{(0)}$ and view weights $V_{\text{global}}^{(0)}$
 Set communication rounds $T$, local iterations $E$, tensor ranks $\{r_{2},r_{3}\}$
for $t=0,1,\ldots,T-1$ do
  Broadcast: Send $\mathcal{G}_{\text{global}}^{(t)},\mathbf{P}_{\text{global}}^{(t)},\mathbf{Q}_{\text{global}}^{(t)},\mathbf{R}_{\text{global}}^{(t)},V_{\text{global}}^{(t)}$ to all clients
  Local Computation (in parallel for all clients):
  for each client $\ell\in\{1,\ldots,M\}$ do
   // Tensorize local multi-view data
   Organize $X_{[\ell]}=\{X_{[\ell]}^{h}\}_{h=1}^{s(\ell)}$ into tensor $\mathcal{X}_{[\ell]}$
   // Personalized tensor factor initialization
   $\mathcal{G}_{[\ell]}^{(0)}\leftarrow\lambda_{\ell}^{core}\mathcal{G}_{\text{global}}^{(t)}+(1-\lambda_{\ell}^{core})\mathcal{G}_{[\ell]}$
   $\mathbf{P}_{[\ell]}^{(0)}\leftarrow\lambda_{\ell}^{P}\mathbf{P}_{\text{global}}^{(t)}+(1-\lambda_{\ell}^{P})\mathbf{P}_{[\ell]}$
   $\mathbf{Q}_{[\ell]}^{(0)}\leftarrow\lambda_{\ell}^{Q}\mathbf{Q}_{\text{global}}^{(t)}+(1-\lambda_{\ell}^{Q})\mathbf{Q}_{\ell}$
   $\mathbf{R}_{[\ell]}^{(0)}\leftarrow\lambda_{\ell}^{R}\mathbf{R}_{\text{global}}^{(t)}+(1-\lambda_{\ell}^{R})\mathbf{R}_{[\ell]}$
   $A_{[\ell]}^{(0)}\leftarrow\rho_{\ell}A_{\text{global}}^{(t)}+(1-\rho_{\ell})A_{[\ell]}$
   $V_{[\ell]}^{(0)}\leftarrow\rho_{\ell}V_{\text{global}}^{(t)}+(1-\rho_{\ell})V_{[\ell]}$
   // Local tensorized heat-kernel enhanced clustering
   for $e=1,2,\ldots,E+1$ do
    Compute $\delta_{[\ell]ij}^{h}$ using Eq. 61
    Reconstruct $\mathcal{A}_{[\ell]}\leftarrow\mathcal{G}_{[\ell]}\times_{1}\mathbf{P}_{[\ell]}\times_{2}\mathbf{Q}_{[\ell]}\times_{3}\mathbf{R}_{[\ell]}$
    Perform Algorithm 2
   end for
   // Compute tensorized aggregation statistics
   Compute $S_{[\ell]}^{core},S_{[\ell]}^{factors},S_{[\ell]}^{views},S_{[\ell]}^{quality}$
   Send statistics to server
  end for
  Server Tensor Aggregation:
@@
  Compute client weights $\omega_{\ell}$ using tensor quality measures
  Update $\mathcal{G}_{\text{global}}^{(t+1)}$ using Eq. 103
  Update $\mathbf{P}_{\text{global}}^{(t+1)}$ using Eq. 104
  Update $\mathbf{Q}_{\text{global}}^{(t+1)}$ using Eq. 105
  Update $\mathbf{R}_{\text{global}}^{(t+1)}$ using Eq. 106
  Update $A_{\text{global}}^{(t+1)}$ using Eq. 90
  Update $V_{\text{global}}^{(t+1)}$ using Eq. 91
  Adaptive Personalization:
  Update tensor-specific personalization parameters $\{\lambda_{\ell}^{core},\lambda_{\ell}^{P},\lambda_{\ell}^{Q},\lambda_{\ell}^{R},\rho_{\ell}\}$
end for
return Global tensor model $(\mathcal{G}_{\text{global}}^{(T)},\mathbf{P}_{\text{global}}^{(T)},\mathbf{Q}_{\text{global}}^{(T)},\mathbf{R}_{\text{global}}^{(T)},A_{\text{global}}^{(T)},V_{\text{global}}^{(T)})$ and personalized models $\{(\mathcal{G}_{[\ell]}^{(T)},\mathbf{P}_{[\ell]}^{(T)},\mathbf{Q}_{[\ell]}^{(T)},\mathbf{R}_{[\ell]}^{(T)},A_{[\ell]}^{(T)},V_{[\ell]}^{(T)})\}_{\ell=1}^{M}$
@@ 6.7.3 Computational Complexity
The local clustering complexity per client $\ell$ per iteration is $\mathcal{O}(n(\ell)c(\ell)s(\ell)D(\ell))$ where $D(\ell)=\sum_{h=1}^{s(\ell)}d_{h}^{(\ell)}$ is the total local dimensionality.
@@ 6.7.4 Tensorized Complexity Analysis
The dominant term is $\mathcal{O}(n(\ell)c(\ell)D(\ell)s(\ell))$ when $r_{2},r_{3}\ll\min(D(\ell),s(\ell))$. ∎
@@
The tensorized federated protocol requires $\mathcal{O}(c(\ell)r_{2}r_{3}+c(\ell)^{2}+\bar{D}(\ell)r_{2}+\bar{s}(\ell)r_{3})$ communication per round, compared to $\mathcal{O}(c(\ell)\bar{D}(\ell)\bar{s}(\ell))$ for the non-tensorized version.
The tensorized approach provides significant communication savings when $r_{2}\ll\bar{D}(\ell)$ and $r_{3}\ll\bar{s}(\ell)$, which is typically the case in practice.

### v3

@@ 4.5.1 Global Objective with Personalization
$J_{\text{global}}=\sum_{\ell=1}^{M}\left[\omega_{\ell}J_{FedHK}^{(\ell)}(U_{[\ell]},A_{[\ell]},V_{[\ell]})+\gamma\mathcal{R}_{person}(A_{[\ell]},A_{\text{global}})+\eta\mathcal{R}_{view}(V_{[\ell]},V_{\text{global}})\right]$ (93)
where $\omega_{\ell},\gamma,\eta$ are hyperparameters controlling the influence of local objectives and regularization terms on the global model. Specifically, $omega_{\ell}$ weights the contribution of client $\ell$’s local objective $J_{FedHK}^{(\ell)}$, while $\gamma,\eta$ control personalization strength and trade offs between local and global models, respectively. $\mathcal{R}_{person}(A_{[\ell]},A_{\text{global}}=\sum_{h,k}\|A_{[\ell]}^{h}-A_{\text{global}}^{h}\|_{F}^{2}$ ensures coordination between local and global centers, and $\mathcal{R}_{view}(V_{[\ell]},V_{\text{global}})=\|V_{[\ell]}-V_{\text{global}}\|_{2}^{2}$ maintains view weight consistency across clients.
@@
4.6 Tensorized Federated Aggregation with Personalization
4.6.1 Tensorized Global Objective with Personalization
$J_{\text{tensor(global)}}=\sum_{\ell=1}^{M}\left[\omega_{\ell}J_{[\ell]}^{tensor}(\mathcal{U}_{[\ell]},\mathcal{G}_{[\ell]},\mathbf{P}_{[\ell]},\mathbf{Q}_{[\ell]},\mathbf{R}_{[\ell]},\mathbf{V}_{[\ell]})+\gamma\mathcal{R}_{tensor}(\mathcal{G}_{[\ell]},\mathbf{P}_{[\ell]},\mathbf{Q}_{[\ell]},\mathbf{R}_{[\ell]})+\eta\mathcal{R}_{view}(V_{[\ell]},V_{\text{global}})\right]$
where the tensorized regularization terms are:
$\mathcal{R}_{tensor}=\|\mathcal{G}_{[\ell]}-\mathcal{G}_{\text{global}}\|_{F}^{2}+\|\mathbf{P}_{[\ell]}-\mathbf{P}_{\text{global}}\|_{F}^{2}+\|\mathbf{Q}_{[\ell]}-\mathbf{Q}_{\text{global}}\|_{F}^{2}+\|\mathbf{R}_{[\ell]}-\mathbf{R}_{\text{global}}\|_{F}^{2}\mathcal{R}_{view}(V_{[\ell]},V_{\text{global}})=\|V_{[\ell]}-V_{\text{global}}\|_{2}^{2}$
@@ 4.6.3 Personalized Update Mechanism
$V_{[\ell]}^{(t+1)}$ $=\rho_{\ell}V_{[\ell]}^{(t)}+(1-\rho_{\ell})V_{\text{global}}^{(t+1)}$ (102)
@@
$\lambda_{\ell}^{(t+1)}=\lambda_{\ell}^{(t)}-\beta\nabla_{\lambda_{\ell}}J_{[\ell]}(\lambda_{\ell}A_{[\ell]}^{(t)}+(1-\lambda_{\ell})A_{\text{global}}^{(t)})$ (103)
@@ 4.6.5 Tensorized Personalized Update Mechanism
4.7 Complete Algorithm for FedHK-MVC-Person with PARAFAC2 Decomposition
@@
Algorithm 4 outlines the complete procedure for the FedHK-MVC-Person framework utilizing PARAFAC2 tensor decomposition. The algorithm encompasses server initialization, local client computations with personalization, and secure aggregation of tensor factors. In order to achieve a personalized FedHK-Tucker, two changes must be made. Initially, adjustments are made to the tensor factor implication stage and its localization tensorized HKC. Finally, the server aggregation phase is modified to accommodate the Tucker decomposition factors.
Algorithm 4 FedHK-MVC-Person with PARAFAC2 Decomposition (Personalized FedHK-PARAFAC2)
0: Multi-view data $\{X_{[\ell]}\}_{\ell=1}^{M}$, clusters $c$, tensor ranks $(r_{1},r_{2},r_{3})$, fuzzifier $m>1$, view exponent $\alpha>1$, rounds $T$, local epochs $E$
0: Global model $\Theta_{\text{global}}^{(T)}$, personalized models $\{\Theta_{[\ell]}^{(T)}\}_{\ell=1}^{M}$
1:/* Phase 1: Server Initialization */
2: Initialize $\mathcal{G}_{\text{global}}^{(0)},\mathbf{P}_{\text{global}}^{(0)},\mathbf{Q}_{\text{global}}^{(0)},\mathbf{R}_{\text{global}}^{(0)},V_{\text{global}}^{(0)}$ randomly
3: Initialize personalization parameters $\{\lambda_{\ell}^{core},\lambda_{\ell}^{P},\lambda_{\ell}^{Q},\lambda_{\ell}^{R},\rho_{\ell}\}_{\ell=1}^{M}\leftarrow 0.5$
4:for $t=0$ to T1T-1 do
5:  /* Phase 2: Server Broadcast */
6:  Broadcast $\Theta_{\text{global}}^{(t)}=\{\mathcal{G}_{\text{global}}^{(t)},\mathbf{P}_{\text{global}}^{(t)},\mathbf{Q}_{\text{global}}^{(t)},\mathbf{R}_{\text{global}}^{(t)},V_{\text{global}}^{(t)}\}$ to all clients
7:  /* Phase 3: Parallel Local Computation */
8:  for each client $\ell\in\{1,\ldots,M\}$ in parallel do
9:   Step 3.1: Data Tensorization
10:   Construct tensor $\mathcal{X}_{[\ell]}\in\mathbb{R}^{n(\ell)\times D_{[\ell]}\times s(\ell)}$ from $\{X_{[\ell]}^{h}\}_{h=1}^{s(\ell)}$ via Eq. 65
11:   Step 3.2: Personalized Initialization
12:   $\mathcal{G}_{[\ell]}^{(0)}\leftarrow\lambda_{\ell}^{core}\mathcal{G}_{\text{global}}^{(t)}+(1-\lambda_{\ell}^{core})\mathcal{G}_{[\ell]}^{prev}$
13:   $\mathbf{P}_{[\ell]}^{(0)},\mathbf{Q}_{[\ell]}^{(0)},\mathbf{R}_{[\ell]}^{(0)}\leftarrow$ analogous personalized blending
14:   $V_{[\ell]}^{(0)}\leftarrow\rho_{\ell}V_{\text{global}}^{(t)}+(1-\rho_{\ell})V_{[\ell]}^{prev}$
15:   Step 3.3: Local Optimization Loop
16:   for $e=1$ to $E$ do
17:    Compute heat-kernel coefficients $\delta_{[\ell]ij}^{h}$ via Eq. 68
18:    Reconstruct centers: $\mathcal{A}_{[\ell]}\leftarrow\mathcal{G}_{[\ell]}\times_{1}\mathbf{P}_{[\ell]}\times_{2}\mathbf{Q}_{[\ell]}\times_{3}\mathbf{R}_{[\ell]}$
19:    Update memberships $U_{[\ell]}^{*}$ via Eq. 75
20:    Update core tensor $\mathcal{G}_{[\ell]}$ via ALS (Eq. 76)
21:    Update factors $\mathbf{P}_{[\ell]},\mathbf{Q}_{[\ell]},\mathbf{R}_{[\ell]}$ via ALS (Eqs. 7981)
22:    Update view weights $V_{[\ell]}$ via Eq. 77
23:   end for
24:   Step 3.4: Compute Privacy-Preserving Statistics
25:   $S_{[\ell]}\leftarrow\{S_{[\ell]}^{core},S_{[\ell]}^{factors},S_{[\ell]}^{views},S_{[\ell]}^{quality},J_{[\ell]}^{tensor},n(\ell)\}$
26:   Send $S_{[\ell]}$ to server
27:  end for
28:  /* Phase 4: Server Aggregation */
@@
29:  Compute client weights: $\omega_{\ell}^{(t)}\leftarrow\frac{\exp(-\tau J_{[\ell]}^{tensor})}{\sum_{\ell^{\prime}}\exp(-\tau J_{[\ell^{\prime}]}^{tensor})}$
30:  Aggregate: $\mathcal{G}_{\text{global}}^{(t+1)},\mathbf{P}_{\text{global}}^{(t+1)},\mathbf{Q}_{\text{global}}^{(t+1)},\mathbf{R}_{\text{global}}^{(t+1)}$ via Eqs. 108111
31:  Aggregate: $V_{\text{global}}^{(t+1)}$ via Eq. 100
32:  /* Phase 5: Adaptive Personalization Update */
33:  for each client $\ell\in\{1,\ldots,M\}$ do
34:   $\lambda_{\ell}^{(\cdot)}\leftarrow\lambda_{\ell}^{(\cdot)}-\beta\nabla_{\lambda_{\ell}^{(\cdot)}}J_{[\ell]}^{tensor}$ {Gradient-based adaptation}
35:   Clip: $\lambda_{\ell}^{(\cdot)}\leftarrow\max(0,\min(1,\lambda_{\ell}^{(\cdot)}))$
36:  end for
37:end for
38:return $\Theta_{\text{global}}^{(T)}$, $\{\Theta_{[\ell]}^{(T)}\}_{\ell=1}^{M}$
@@ 4.8.3 Computational Complexity
The local clustering complexity per client $\ell$ per iteration is $\mathcal{O}(n(\ell)c(\ell)s(\ell)D(\ell))$ where $D(\ell)=\sum_{h=1}^{s(\ell)}d_{h}^{(\ell)}$ is the total local dimensionality across all views.
Proof.
The computational complexity of the E-FKMVC algorithm at each client $\ell$ can be analyzed by examining the dominant operations performed during a single iteration. The algorithm consists of three main update steps, each contributing to the overall time complexity:
Step 1: Membership Matrix Update. Computing the membership values $\mu_{[\ell]ik}^{*}$ using Eq. 29 requires evaluating the FKED distance between each data point $i$ and each cluster center $k$ across all views $h$. For each of the $n(\ell)$ data points and $c(\ell)$ clusters, we must sum over $s(\ell)$ views, with each view contributing $d_{h}^{(\ell)}$ feature-wise distance computations. This yields a complexity of $\mathcal{O}(n(\ell)\cdot c(\ell)\cdot\sum_{h=1}^{s(\ell)}d_{h}^{(\ell)})=\mathcal{O}(n(\ell)c(\ell)D(\ell))$.
Step 2: Cluster Center Update. Updating each cluster center $a_{[\ell]kj}^{h}$ using Eq. 30 requires computing a weighted sum over all $n(\ell)$ data points for each of the $c(\ell)$ clusters, across all features in all views. This contributes $\mathcal{O}(n(\ell)c(\ell)D(\ell))$ operations.
Step 3: View Weight Update. Computing the view weights $v_{[\ell]h}$ using Eq. 31 requires summing distance contributions over all data points, clusters, and views, resulting in $\mathcal{O}(n(\ell)c(\ell)s(\ell))$ operations.
Combining these contributions, the dominant term is $\mathcal{O}(n(\ell)c(\ell)D(\ell))$. Since $D(\ell)=\sum_{h=1}^{s(\ell)}d_{h}^{(\ell)}\geq s(\ell)$ in practical scenarios where each view has at least one feature, we can express the total complexity as $\mathcal{O}(n(\ell)c(\ell)s(\ell)D(\ell))$ to explicitly highlight the dependence on all key parameters. ∎
Practical Interpretation: The time complexity result has important practical implications for deploying the E-FKMVC algorithm in real-world federated learning systems:
  • Linear scaling with data size: The algorithm scales linearly with the number of local data points $n(\ell)$, making it suitable for clients with varying dataset sizes.
  • Linear scaling with cluster count: The complexity grows linearly with the number of clusters $c(\ell)$, enabling efficient computation even when modeling fine-grained cluster structures.
  • Multi-view overhead: The dependence on $s(\ell)$ and $D(\ell)$ reflects the computational cost of integrating information across multiple data views. This overhead is a necessary trade-off for achieving improved clustering accuracy through multi-view learning.
  • Comparison with single-view clustering: Standard fuzzy c-means clustering has complexity $\mathcal{O}(n\cdot c\cdot d)$ for a single view. Our multi-view extension introduces a multiplicative factor proportional to the number of views, which is typical for multi-view learning algorithms that must process and integrate information from each view separately.
@@ 4.8.4 Tensorized Complexity Analysis
The dominant term is $\mathcal{O}(n(\ell)c(\ell)D(\ell)s(\ell))$ when $r_{2},r_{3}\ll\min(D(\ell),s(\ell))$. Overall, the complexity is $\mathcal{O}(n(\ell)c(\ell)r_{2}r_{3}+n(\ell)cD(\ell)s(\ell))$. While the tensorized approach introduces additional overhead from tensor operations, it remains efficient for low-rank decompositions. Table 5 summarizes the complexity comparison Table 5 summarizes the complexity comparison between standard (E-FKMVC) and tensorized federated clustering (FedHK-PARAFAC2, FedHK-Tucker, and FedHK-MVC-Person). ∎
Table 5: Complexity Comparison of Federated Clustering Methods
Method Time Complexity Communication Complexity E-FKMVC $\mathcal{O}(n(\ell)c(\ell)s(\ell)D(\ell))$ $\mathcal{O}(c(\ell)\bar{d}_{h}^{(\ell)}\bar{s}(\ell))$ FedHK-PARAFAC2 $\mathcal{O}(n(\ell)c(\ell)r_{2}r_{3}+n(\ell)cD(\ell)s(\ell))$ $\mathcal{O}(c(\ell)r_{2}r_{3}+c(\ell)^{2}+\bar{D}(\ell)r_{2}+\bar{s}(\ell)r_{3})$ FedHK-Tucker $\mathcal{O}(n(\ell)c(\ell)r_{2}r_{3}+n(\ell)cD(\ell)s(\ell))$ $\mathcal{O}(c(\ell)r_{2}r_{3}+c(\ell)^{2}+\bar{D}(\ell)r_{2}+\bar{s}(\ell)r_{3})$ FedHK-MVC-Person $\mathcal{O}(n(\ell)c(\ell)r_{2}r_{3}+n(\ell)cD(\ell)s(\ell))$ $\mathcal{O}(c(\ell)r_{2}r_{3}+c(\ell)^{2}+\bar{D}(\ell)r_{2}+\bar{s}(\ell)r_{3})$
As can be observed in Table 5, the tensorized methods introduce additional computational overhead due to tensor operations, but they remain efficient for low-rank decompositions. While FedHK-PARAFAC2, FedHK-Tucker, and FedHK-MVC-Person share the same asymptotic time complexity, their communication complexities are also identical due to the similar structure of shared statistics.
@@
The tensorized federated protocol requires $\mathcal{O}(c(\ell)r_{2}r_{3}+c(\ell)^{2}+\bar{D}(\ell)r_{2}+\bar{s}(\ell)r_{3})$ communication per round, compared to $\mathcal{O}(c(\ell)\bar{D}(\ell)\bar{s}(\ell))$ for the non-tensorized version. The communication savings are significant when $r_{2}\ll\bar{D}(\ell)$ and $r_{3}\ll\bar{s}(\ell)$.
Lemma 1 (Tensorized Communication Complexity).
The communication complexity per client $\ell$ per round for the tensorized federated protocol is $\mathcal{O}(c(\ell)r_{2}r_{3}+c(\ell)^{2}+\bar{D}(\ell)r_{2}+\bar{s}(\ell)r_{3})$.
Proof.
The communication complexity arises from the size of the shared statistics:
  • Core tensor statistics: $\mathcal{O}(c(\ell)r_{2}r_{3})$. This accounts for the shared core tensor slices for each cluster.
  • Factor matrices statistics: $\mathcal{O}(c(\ell)^{2}+\bar{D}(\ell)r_{2}+\bar{s}(\ell)r_{3})$. This includes the shared factor matrices $\mathbf{P}_{[\ell]},\mathbf{Q}_{[\ell]},\mathbf{R}_{[\ell]}$.
  • View weights statistics: $\mathcal{O}(s(\ell))$. This accounts for the shared view importance weights.
  • Quality statistics: $\mathcal{O}(1)$. This includes the shared clustering quality measure and data size.
Summing these contributions yields the total communication complexity of $\mathcal{O}(c(\ell)r_{2}r_{3}+c(\ell)^{2}+\bar{D}(\ell)r_{2}+\bar{s}(\ell)r_{3})$. This is significantly lower than the non-tensorized complexity of $\mathcal{O}(c(\ell)\bar{D}(\ell)\bar{s}(\ell))$ when the tensor ranks $r_{2},r_{3}$ are much smaller than the average feature dimensions $\bar{D}(\ell)$ and number of views $\bar{s}(\ell)$. ∎
@@
Tensorized approximations enable efficient representation of complex multi-view cluster structures with controlled error. While tensorized communication introduces additional overhead from tensor operations, it remains efficient for low-rank decompositions.